### v1

# Optimized Many-Hypercube Codes toward Lower Logical Error Rates and Earlier Realization

Abstract
Many-hypercube codes [H. Goto, Sci. Adv. 10, eadp6388 (2024)], concatenated [[n,n2,2]]{[[n,n-2,2]]} quantum error-detecting codes ($n$ is even), have recently been proposed as high-rate quantum codes suitable for fault-tolerant quantum computing. However, the original many-hypercube codes with n=6{n=6} have large code block sizes at high concatenation levels (216 and 1296 physical qubits per block at levels 3 and 4, respectively), making not only experimental realization difficult but also logical error rates high. Toward earlier experimental realization and lower logical error rates, here we investigate smaller many-hypercube codes obtained by concatenating $[[6,4,2]]$ and/or $[[4,2,2]]$ codes, where, e.g., $D_{6,4,4}$ denotes the many-hypercube code using $[[6,4,2]]$ at level 1 and $[[4,2,2]]$ at levels 2 and 3. As a result, we found a surprising fact: $D_{6,4,4}$ ($D_{6,6,4,4}$) can achieve lower block error rates than $D_{4,4,4}$ ($D_{4,4,4,4}$), despite its higher encoding rate. Focusing on level 3, we also developed efficient fault-tolerant encoders realizing about 60% overhead reduction while maintaining or even improving the performance, compared to the original design. Using them, we numerically confirmed that $D_{6,4,4}$ also achieves the best performance for logical controlled-NOT gates in a circuit-level noise model. These results will be useful for early experimental realization of fault-tolerant quantum computing with high-rate quantum codes.
I Introduction
Building a reliable quantum computer is challenging, because of noises, e.g., due to decoherence and imperfection of quantum-system control. To overcome this difficulty, quantum error-correcting codes [1, 2, 3, 4, 5, 6, 7] and fault-tolerant quantum computation (FTQC) with them [8, 9, 10] have been proposed. Since the threshold theorem that arbitrarily long quantum computations can be performed reliably if physical error rates are lower than a threshold was proved by using concatenated quantum codes [11], FTQC had been based on concatenated quantum codes [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. This approach culminated by Knill’s $C_{4}/C_{6}$} scheme showing the threshold over 1%, which is based on the concatenation of simple quantum error-detecting codes [16]. At that time, however, it was extremely difficult to experimentally realize these proposals, in particular, nonlocal gates such as transversal controlled-NOT (CNOT) gates [10].
Topological codes [24], in particular, the surface code [25, 26] have become more popular since it was shown that the surface code can achieve comparable thresholds only using local operations [27, 28, 29, 30, 31]. Moreover, a compact version called the rotated surface code and a simple two-qubit gate approach called lattice surgery were also proposed [32]. The surface-code approach to FTQC is particularly suitable for superconducting-circuit implementations, leading to recent experimental demonstrations of the surface code with superconducting circuits [33, 34, 35, 36].
The (rotated) surface code encodes a single logical qubit into $d^{2}$ physical qubits, where $d$ is the code distance. Thus, the encoding rate $R=1/d^{2}$ becomes low for large code block sizes, resulting in large resource overheads. To address this overhead issue, quantum low-density parity-check (qLDPC) codes [43, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47], which can achieve a constant encoding rate even for large code distances, have recently attracted much attention [48, 49, 50, 51, 52, 53]. This expectation is enhanced by recent experimental advances, in particular, in ion-trap [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67] and neutral-atom [68, 69, 70, 71, 72, 73, 74, 75, 76, 77] platforms, because long-range interactions beyond nearest-neighbor ones necessary for high-rate codes such as qLDPC codes [78, 79] are experimentally feasible by moving qubits in these platforms. However, parallel execution of addressable logical gates with low overheads is still challenging for qLDPC codes, the study on which is ongoing [80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102].
Alternative high-rate codes, namely, non-LDPC high-rate quantum codes have also been proposed toward low-overhead FTQC [103, 104, 105, 106, 107, 108, 109, 110]. Among them, concatenated high-rate quantum codes are particularly promising in the sense that fault-tolerant logical-state preparation can be achieved by level-by-level encoding [16, 22, 64, 103, 104, 105]. In this work, we focus on concatenated [[n,n2,2]]{[[n,n-2,2]]} quantum error-detecting codes ($n$ is even), which are called many-hypercube (MHC) codes because the code structure can be described by using many $L$-dimensional hypercubes ($L$ is the concatenation level) corresponding to logical qubits (logical Pauli operators) [104, 107, 109]. Note that the error-detecting codes, which are also called iceberg codes [60], are the simplest, highest-rate Calderbank-Shor-Steane (CSS) codes [3, 4, 5]. In the original work [104], concatenated $[[6,4,2]]$ codes were adopted because of the good balance between the high rate and code block size of $[[6,4,2]]$. Recently, smaller MHC codes with $[[4,2,2]]$ have been studied [107, 109]. Remarkably, the smallest level-2 MHC code $4^{2},2^{2},2^{2}]]$ has been realized experimentally using neutral atoms [77]. Smaller block sizes will lead to not only earlier experimental realization, but also lower logical error rates, though the encoding rates become lower. For this reason, in this work we investigate the MHC codes with $[[6,4,2]]$ and/or $[[4,2,2]]$, where the MHC code with $n_{L},n_{L}-2,2]]$} at level $L$ is denoted by $D_{n_{1},n_{2},\ldots}$. As a result, we found that $D_{6,4,4}$ ($D_{6,6,4,4}$) can achieve lower block error rates than $D_{4,4,4}$ ($D_{4,4,4,4}$), despite its higher encoding rate. In addition, it turned out that $D_{4,4,6,6}$ is the worst at level 4. This is a notable result, because in a previous study [107], it was assumed that we should use smaller codes at lower levels, and consequently $D_{4,4,6,6}$ was adopted in the previous study. We also develop efficient fault-tolerant encoders for level-3 MHC codes, which can reduce overheads by about 60% compared to the original design [104]. Using them, we confirmed that $D_{6,4,4}$ can also achieve the highest performance for logical CNOT gates at level 3.
This paper is organized as follows. In Sec. II, we explain the definitions of the MHC codes in detail using figures visualizing their code structures. In Sec. III, we show numerical results for block error rates using a bit-flip (depolarizing) error model. In Sec. IV, we present our new encoders for level-3 MHC codes comparing them with the original ones. In Sec. V, we show numerical results for logical-CNOT error rates at level 3 using the new and original encoders. Finally, we summarize the results and discuss the outlook in Sec. VI.
II Definitions of MHC codes
Here we present the detailed definitions of the MHC codes from level 1 to level 4. We will use figures to show the code structures, but note that they are conceptual and we do not have to realize such qubit configurations in real space.
II.1 $D_{4}$ and $D_{6}$
In this work, we use ${[[4,2,2]]}~D_{4}$ and/or ${[[6,4,2]]}~D_{6}$ as the base codes for the MHC codes. Both have only two stabilizers, all $Z$ and all $X$, and their logical Pauli operators have weight of 2, as shown explicitly below, leading to code distance of 2.
The logical Pauli operators and stabilizers for $D_{4}$ are defined as follows:
$Z_{1}^{(1)}$ $=Z_{1}Z_{2},$ $X_{1}^{(1)}$ $=X_{2}X_{3},$
$Z_{2}^{(1)}$ $=Z_{2}Z_{3},$ $X_{2}^{(1)}$ $=X_{1}X_{2},$
$SZ^{(1)}$ $=Z_{1}Z_{2}Z_{3}Z_{4},$ $SX^{(1)}$ $=X_{1}X_{2}X_{3}X_{4},$
where $Z_{i}$ and $X_{i}$ are the Pauli operators for the $i$th physical qubit, which is denoted by $\mathrm{Q}_{i}$, and the superscript suggests that they are the operators used at level 1 of MHC codes. By aligning the four physical qubits along the virtual $x$ axis, where $\mathrm{Q}_{i}$ is placed at $x=i$, the above operators can be visualized as shown in Fig. 1(a).
The logical Pauli operators and stabilizers for $D_{6}$ are defined as follows:
$Z_{1}^{(1)}$ $=Z_{1}Z_{2},$ $X_{1}^{(1)}$ $=X_{2}X_{3},$
$Z_{2}^{(1)}$ $=Z_{2}Z_{3},$ $X_{2}^{(1)}$ $=X_{1}X_{2},$
$Z_{3}^{(1)}$ $=Z_{4}Z_{5},$ $X_{3}^{(1)}$ $=X_{5}X_{6},$
$Z_{4}^{(1)}$ $=Z_{5}Z_{6},$ $X_{4}^{(1)}$ $=X_{4}X_{5},$
$SZ^{(1)}$ $=Z_{1}Z_{2}Z_{3}Z_{4}Z_{5}Z_{6},$ $SX^{(1)}$ $=X_{1}X_{2}X_{3}X_{4}X_{5}X_{6}.$
Similarly to the case of $D_{4}$, these can be visualized as shown in Fig. 1(b). There are different definitions of [[n,n2,2]][[n,n-2,2]] codes [60], we adopt the above definitions because of the simple visualization [104].
Figure 1: Definitions of (a) $D_{4}$ and (b) $D_{6}$. The circles represent physical qubits. The top two show logical Pauli operators and the bottom two show stabilizers. $Z$ and $X$ operators are shown in blue and red, respectively. Solid and dashed line segments connecting qubits are guides to the eye.
II.2 Level-2 MHC codes
A level-2 MHC code $D_{4,4}$ is obtained by concatenating $D_{4}$ with itself as follows. We first prepare four level-1 code blocks encoded with $D_{4}$. Referring to $Z_{i}^{(1)}$ and $X_{i}^{(1)}$ of the $j$th block as $Z_{i,j}^{(1)}$ and $X_{i,j}^{(1)}$, respectively, the logical Pauli operators of $D_{4,4}$ are defined as
$Z_{i,1}^{(2)}$ $=Z_{i,1}^{(1)}Z_{i,2}^{(1)},X_{i,1}^{(2)}=X_{i,2}^{(1)}X_{i,3}^{(1)},$
$Z_{i,2}^{(2)}$ $=Z_{i,2}^{(1)}Z_{i,3}^{(1)},X_{i,2}^{(2)}=X_{i,1}^{(1)}X_{i,2}^{(1)}.$
Note that all of them have weight of $2^{2}=4$}, leading to code distance of 4. Also, there are $2^{2}=4$} logical qubits. Thus, $D_{4,4}$ is a $[[16,4,4]]$ code. In addition to the original stabilizers of $D_{4}$, the following stabilizers are also introduced:
$SZ_{i}^{(2)}$ $=Z_{i,1}^{(1)}Z_{i,2}^{(1)}Z_{i,3}^{(1)}Z_{i,4}^{(1)},SX_{i}^{(2)}=X_{i,1}^{(1)}X_{i,2}^{(1)}X_{i,3}^{(1)}X_{i,4}^{(1)}.$
By stacking the four level-1 code blocks along the virtual $y$ axis, where $\mathrm{Q}_{i}$ of the $j$th block, which is denoted by $\mathrm{Q}_{i,j}$, is placed at (x,y)=(i,j){(x,y)=(i,j)}, the above operators can be visualized as shown in Fig. 2(a). Note that each logical Pauli operator corresponds to a square (two-dimensional hypercube).
Similarly, another level-2 MHC code $D_{6,4}$ obtained by using $D_{6}$ at level 1 and $D_{4}$ at level 2 is defined by the same equations as $D_{4,4}$ except for $i=1,\ldots,6$} in the case of $D_{6,4}$. $D_{6,4}$ is a $6\cdot 4,4\cdot 2,2^{2}]]=[[24,8,4]]$ code. The operators for $D_{6,4}$ are visualized in Fig. 2(b).
Another level-2 MHC code $D_{4,6}$ obtained by using $D_{4}$ at level 1 and $D_{6}$ at level 2 is defined as follows:
$Z_{i,1}^{(2)}$ $=Z_{i,1}^{(1)}Z_{i,2}^{(1)},X_{i,1}^{(2)}=X_{i,2}^{(1)}X_{i,3}^{(1)},$
$Z_{i,2}^{(2)}$ $=Z_{i,2}^{(1)}Z_{i,3}^{(1)},X_{i,2}^{(2)}=X_{i,1}^{(1)}X_{i,2}^{(1)},$
$Z_{i,3}^{(2)}$ $=Z_{i,4}^{(1)}Z_{i,5}^{(1)},X_{i,3}^{(2)}=X_{i,5}^{(1)}X_{i,6}^{(1)},$
$Z_{i,4}^{(2)}$ $=Z_{i,5}^{(1)}Z_{i,6}^{(1)},X_{i,4}^{(2)}=X_{i,4}^{(1)}X_{i,5}^{(1)},$
$SZ_{i}^{(2)}$ $=Z_{i,1}^{(1)}Z_{i,2}^{(1)}Z_{i,3}^{(1)}Z_{i,4}^{(1)}Z_{i,5}^{(1)}Z_{i,6}^{(1)},$
$SX_{i}^{(2)}$ $=X_{i,1}^{(1)}X_{i,2}^{(1)}X_{i,3}^{(1)}X_{i,4}^{(1)}X_{i,5}^{(1)}X_{i,6}^{(1)}.$
$D_{6,6}$ is defined similarly. $D_{4,6}$ and $D_{6,6}$ are $4\cdot 6,2\cdot 4,2^{2}]]=[[24,8,4]]$ and $6^{2},4^{2},2^{2}]]=[[36,16,4]]$ codes, respectively. These operators are visualized in Figs. 2(c) and 2(d), respectively.
Figure 2: Visualization of level-2 MHC codes. The circles represent physical qubits. Gray squares correspond to logical Pauli operators. In each figure, the top shows logical Pauli operators for $\mathrm{Q}_{1,1}$ and the bottom shows examples of stabilizers by highlighting them in blue and red for $Z$ and $X$ operators, respectively.
II.3 Level-3 MHC codes
A level-3 MHC code $D_{4,4,4}$ using four $D_{4,4}$ blocks is defined as follows:
$Z_{i,j,1}^{(3)}$ $=Z_{i,j,1}^{(2)}Z_{i,j,2}^{(2)},X_{i,j,1}^{(3)}=X_{i,j,2}^{(2)}X_{i,j,3}^{(2)},$
$Z_{i,j,2}^{(3)}$ $=Z_{i,j,2}^{(2)}Z_{i,j,3}^{(2)},X_{i,j,2}^{(3)}=X_{i,j,1}^{(2)}X_{i,j,2}^{(2)}.$
$SZ_{i,j}^{(3)}$ $=Z_{i,j,1}^{(2)}Z_{i,j,2}^{(2)}Z_{i,j,3}^{(2)}Z_{i,j,4}^{(2)},$
$SX_{i,j}^{(3)}$ $=X_{i,j,1}^{(2)}X_{i,j,2}^{(2)}X_{i,j,3}^{(2)}X_{i,j,4}^{(2)},$
where $Z_{i,j,k}^{(2)}$ and $X_{i,j,k}^{(2)}$ denote $Z_{i,j}^{(2)}$ and $X_{i,j}^{(2)}$, respectively, of the $k$th $D_{4,4}$ block. $D_{4,4,4}$ also has the original stabilizers of $D_{4,4}$. By stacking the four $D_{4,4}$ blocks along the virtual $z$ axis, where $\mathrm{Q}_{i,j}$ of the $k$th block, which is denoted by $\mathrm{Q}_{i,j,k}$, is placed at (x,y,z)=(i,j,k){(x,y,z)=(i,j,k)}, the above operators can be visualized as shown in Fig. 3(a). Each logical Pauli operator corresponds to a cube (three-dimensional hypercube). The other level-3 MHC codes $D_{n_{1},n_{2},4}$ are defined similarly. The code structures of $D_{6,4,4}$ and $D_{6,6,4}$ are visualized in Figs. 3(b) and 3(c), respectively.
Another level-6 MHC code $D_{6,6,6}$ obtained by concatenating $D_{6}$ using six $D_{6,6}$ blocks is defined as follows:
$Z_{i,j,1}^{(3)}$ $=Z_{i,j,1}^{(2)}Z_{i,j,2}^{(2)},X_{i,j,1}^{(3)}=X_{i,j,2}^{(2)}X_{i,j,3}^{(2)},$
$Z_{i,j,2}^{(3)}$ $=Z_{i,j,2}^{(2)}Z_{i,j,3}^{(2)},X_{i,j,2}^{(3)}=X_{i,j,1}^{(2)}X_{i,j,2}^{(2)}.$
$Z_{i,j,3}^{(3)}$ $=Z_{i,j,4}^{(2)}Z_{i,j,5}^{(2)},X_{i,j,3}^{(3)}=X_{i,j,5}^{(2)}X_{i,j,6}^{(2)}.$
$Z_{i,j,4}^{(3)}$ $=Z_{i,j,5}^{(2)}Z_{i,j,6}^{(2)},X_{i,j,4}^{(3)}=X_{i,j,4}^{(2)}X_{i,j,5}^{(2)}.$
$SZ_{i,j}^{(3)}$ $=Z_{i,j,1}^{(2)}Z_{i,j,2}^{(2)}Z_{i,j,3}^{(2)}Z_{i,j,4}^{(2)}Z_{i,j,5}^{(2)}Z_{i,j,6}^{(2)},$
$SX_{i,j}^{(3)}$ $=X_{i,j,1}^{(2)}X_{i,j,2}^{(2)}X_{i,j,3}^{(2)}X_{i,j,4}^{(2)}X_{i,j,5}^{(2)}X_{i,j,6}^{(2)}.$
Some of these operators are shown in Fig. 3(d). The other level-3 MHC codes $D_{n_{1},n_{2},6}$ are defined similarly.
II.4 Level-4 MHC codes
The level-4 MHC codes $D_{n_{1},n_{2},n_{3},4}$ and $D_{n_{1},n_{2},n_{3},6}$ are defined by similar equations to those for $D_{n_{1},n_{2},4}$ and $D_{n_{1},n_{2},6}$, respectively, in the last sectoin. By introducing the fourth virtual axis denoted by $t$ and stacking level-3 $D_{n_{1},n_{2},n_{3}}$ blocks along the $t$ axis, each logical Pauli operator of the level-4 MHC codes corresponds to a four-dimensional hypercube, where $\mathrm{Q}_{i,j,k}$ of the $l$th $D_{n_{1},n_{2},n_{3}}$ block is placed at $(x,y,z,t)=(i,j,k,l)$. This is the origin of the name of MHC codes.
Figure 3: Visualization of level-3 MHC codes. The circles represent physical qubits. Gray cubes correspond to logical Pauli operators. In each figure, the left shows logical Pauli operators for $\mathrm{Q}_{1,1,1}$ and the right shows examples of stabilizers by highlighting them in blue and red for $Z$ and $X$ operators, respectively.
III Performance comparison of MHC codes
Here we numerically evaluate and compare the performance of MHC codes for bit-flip errors. Note that because of the code symmetry with respect to $Z$ and $X$, the performance for bit-flip errors with error probability of $p_{\mathrm{flip}}$ is the same as that for depolarizing errors with error probability of $p_{D}=1.5p_{\mathrm{flip}}$. This is because in the depolarizing error model, one of the three Pauli errors, $X$, $Z$, and Y=iXZ{Y=iXZ}, occurs with the same probability $p_{D}/3$ and therefore a bit-flip ($X$) error occurs with probability of $2p_{D}/3$.
In this numerical simulation, we first prepare an error-free all-zero state encoded with a MHC code. The all-zero state encoded with $D_{4}$ is a four-qubit Greenberger-Home-Zeilinger (GHZ) state,
$|00\rangle_{\mathrm{L}1}=\frac{|0000\rangle+|1111\rangle}{\sqrt{2}},$ (1)
and that with $D_{6}$ is a six-qubit GHZ state,
$|0000\rangle_{\mathrm{L}1}=\frac{|000000\rangle+|111111\rangle}{\sqrt{2}}.$ (2)
Thus, the initial state can be prepared using the encoders in Figs. 5(a) and 5(b), respectively, where we assume no errors on all the operations. Next, we induce bit-flip errors on the physical qubits, then measure all the physical qubits in the $Z$ basis, and finally decode the measurement results using the level-by-level minimum-distance decoder proposed in the original work [104, 111]. Unless all the logical values obtained by the decoding are 0, we consider that the decoding fails, leading to the block error probability. In this work, we used a Python package called Stim [115] for quantum circuit simulations.
Figure 4: All-zero state encoders for (a) $D_{4}$ and (b) $D_{6}$.
Figure 5(a) shows the results for the four level-2 MHC codes. The results are natural in the sense that lower-rate, smaller-size codes can achieve higher performance (lower decoding error probabilities) than higher-rate, larger-size codes. On the other hand, it is noteworthy that $D_{6,4}$ outperforms $D_{4,6}$. This is notable because it has been assumed that smaller-size codes should be used at lower levels [107].
We also evaluated the performance of level-3 and level-4 MHC codes at $p_{\mathrm{flip}}=1\%$} and 2%, respectively, the results of which are shown in Figs. 5(b) and 5(c), respectively. As in the case of level 2, our result clearly suggests that we should use larger-size codes for lower levels for MHC codes, contrary to the previous assumption [107]. More importantly, Figs. 5(b) and 5(c) provide another counterintuitive result that the highest-performance MHC code is not the smallest-size one. The best MHC code is $D_{6,4,4}$ at level 3 and $D_{6,6,4,4}$ at level 4. (The level-4 MHC code chosen in a previous work [107], $D_{4,4,6,6}$, is the worst.) This is the main result of this work.
Figure 5: Performance comparison of MHC codes. (a–c) Decoding error probabilities at level 2–4, respectively, for bit-flip errors.
IV Proposed encoders of level-3 MHC codes
Before logical-CNOT performance evaluation in a circuit-level noise model, here we propose new fault-tolerant encoders for the level-3 MHC codes.
At level 1, we use the standard method for GHZ states with an ancilla physical qubit, as in the original work [104]. Thus, the total number of physical qubits necessary for preparing a $D_{n_{1}}$ all-zero state is $N^{\prime(1)}_{n_{1}}=n_{1}+1$.
At level 2, the encoder in Fig. 6(a) was proposed in the original work [104], where the error-detection gadgets are implemented by the flag method [112] as in Figs. 6(b) and 6(c) for $D_{4}$ and as in Figs. 6(d) and 6(e) for $D_{6}$, and also logical $X$ errors are removed by using a logical ancilla block. The proposed encoder is shown in Fig. 6(f), where the $Z$-error and $X$-error detection is performed at once more efficiently by the method in Figs. 6(g) for $D_{4}$ or 6(h) for $D_{6}$, and also we do not use a logical ancilla block. The simultaneous measurements of weight-4 $Z$ and $X$ stabilizers in Fig. 6(g) were proposed in Ref. 113, where the ancilla qubits play both the roles of error detection and flag, and therefore additional flag qubits are unnecessary. That is, this is more efficient than the methods in Figs. 6(b) and 6(c). The simultaneous measurements of weight-6 $Z$ and $X$ stabilizers in Fig. 6(h) are based on the proposal in Ref. 60, but we modify it by adding a flag qubit to remove correlated $X$ errors, leading to higher performance. To remove logical $X$ errors, instead of using a logical ancilla block, we directly measure logical $Z$ operators of the level-2 MHC codes using physical ancilla qubits, as shown in Fig. 7. Crucial correlated $Z$ errors caused by these measurements can be removed by the following error detection, making the encoder fault-tolerant. To achieve this, the order of the physical CNOT gates in Fig. 7 is important.
The total number of physical qubits necessary for preparing a $D_{n_{1},n_{2}}$ all-zero state is
$N^{\prime(2)}_{n_{1},n_{2}}$ $=(n_{2}+1)N^{(1)}_{n_{1}}+4$
$=n_{2}n_{1}+n_{2}+n_{1}+5$ (3)
for the original encoder and
$N^{\prime(2)}_{n_{1},n_{2}}$ $=n_{2}N^{(1)}_{n_{1}}+(n_{1}-2)+n_{1}/2$
$=n_{2}n_{1}+n_{2}+3n_{1}/2-2$ (4)
for the proposed encoder. (Note that the latter formula holds only for $n_{1},n_{2}=4,6$}.) Thus, the proposed encoder can reduce the overhead by $7-n_{1}/2$.
Figure 6: Fault-tolerant encoders for level-2 MHC codes. Here we show the encoders for only $D_{n,4}$ (n=4,6{n=4,6}), but $D_{n,6}$ encoders are similar. $K^{(1)}$ denotes the number of logical qubits in the level-1 block. (a) Original encoder. EDZL1 and EDXL1 are level-1 $Z$-error and $X$-error detection gadgets, respectively, which are implemented as in (b) and (c), respectively, for $D_{4}$ and as in (d) and (e), respectively, for $D_{6}$. (b,c) $Z$-error (b) and $X$-error (c) detection for $D_{4}$ with a flag qubit. (d,e) $Z$-error (d) and $X$-error (e) detection for $D_{6}$ with a flag qubit. (f) Proposed encoder. LZ is a logical $Z$-error detection gadget implemented as in Fig. 7. EDZXL1 is a level-1 simultaneous $Z$-error and $X$-error detection gadget implemented as in (g) for $D_{4}$ and as in (h) for $D_{6}$. (g,h) Simultaneous $Z$-error and $X$-error detection for $D_{4}$ (g) and $D_{6}$ (h).
Figure 7: Direct measurements of level-2 logical $Z$ operators. Here we show only the cases of $D_{i,4}$ (i=4,6{i=4,6}), but the cases of $D_{i,6}$ are similar. The squares are physical ancilla qubits used for the measurements, which are initially set in $|0\rangle$. (a,b) Gate sequence for the logical $Z$ measurements used in the $D_{4,4}$ and $D_{6,4}$ encoders, respectively. Note that the bottom two level-1 blocks correspond to the top two ones in Fig. 6(f).
The original level-3 encoder shown in Fig. 8(a) is similar to the level-2 encoder in Fig. 6(a), but different implementations based on the Steane method [114] [Figs. 8(b) and 8(c)] are adopted for the error-detection gadgets [104]. The proposed encoder shown in Fig. 8(d) is similar to the level-2 one in Fig. 6(f), but it does not detect logical $X$ errors. This is enough if we use the level-3 logical qubits for quantum computation (not use for level-4 logical state preparation). Also, we perform simultaneous $Z$-error and $X$-error detection for not only level-1 blocks using the method in Fig. 6(g) or 6(h), but also a level-2 block using the method in Fig. 8(e) or 8(f).
The total number of physical qubits necessary for preparing a $D_{n_{1},n_{2},n_{3}}$ all-zero state is
$N^{\prime(3)}_{n_{1},n_{2},n_{3}}$ $=(n_{3}+1)N^{(2)}_{n_{1},n_{2}}+2N^{(2)}_{n_{1},n_{2}}$
$=n_{3}n_{2}n_{1}+n_{3}n_{2}+n_{3}n_{1}+3n_{2}n_{1}$
$+5n_{3}+3n_{2}+3n_{1}+15$ (5)
for the original encoder and
$N^{\prime(3)}_{n_{1},n_{2},n_{3}}$ $=n_{3}N^{(2)}_{n_{1},n_{2}}+(n_{2}/2)N^{(1)}_{n_{1}}+n_{2}(n_{1}/2)$
$=n_{3}n_{2}n_{1}+n_{3}n_{2}+3n_{3}n_{1}/2+n_{2}n_{1}$
$-2n_{3}+n_{2}/2$ (6)
for the proposed encoder. (Note that the latter formula holds only for $n_{1},n_{2},n_{3}=4,6$}.) Thus, the proposed encoder can reduce the overhead by $2n_{2}n_{1}+7n_{3}+5n_{2}/2+3n_{1}+15-n_{3}n_{1}/2$. This is about 60% of the original overhead for all the eight level-3 MHC codes. Thus, the proposed encoder realizes about 60% reduction of overhead.
Figure 8: Fault-tolerant encoders for level-3 MHC codes. Here we show the encoders for only $D_{n_{1},n_{2},4}$ ($n_{1},n_{2}=4,6$}), but $D_{n_{1},n_{2},6}$ encoders are similar. $K^{(L)}$ ($L=1,2$) denotes the number of logical qubits in the level-$L$ block. (a) Original encoder. EDZL2 and EDXL2 are level-2 $Z$-error and $X$-error detection gadgets, respectively, which are implemented as in (b) and (c), respectively. (b,c) $Z$-error (b) and $X$-error (c) detection based on the Steane method. (d) Proposed encoder. EDZXL2 is a level-2 simultaneous $Z$-error and $X$-error detection gadget implemented as in (e) for $D_{n_{1},4}$ and as in (f) for $D_{n_{1},6}$ ($n_{1}=4,6$). EDZXL1 is level-1 simultaneous $Z$-error and $X$-error detection gadgets implemented as in Fig. 6(g) for $D_{4}$ and as in Fig. 6(h) for $D_{6}$. (e,f) Simultaneous $Z$-error and $X$-error detection for $D_{n_{1},4}$ (e) and $D_{n_{1},6}$ (f) ($n_{1}=4,6$).
V Logical CNOT performance
Using the encoders presented in the last section, we evaluated logical CNOT performance for four promising level-3 MHC codes, namely, $D_{4,4,4}$, $D_{6,4,4}$, $D_{6,6,4}$, and $D_{6,6,6}$, in a circuit-level noise model with error rate $p_{\mathrm{circ}}$. As in the original work [104], here we assume that each of physical-qubit zero-state preparation and $Z$-basis measurement is accompanied by a bit flip with probability $p_{\mathrm{circ}}$, each physical CNOT gate is followed by one of 15 two-qubit Pauli errors with equal probability $p_{\mathrm{circ}}/15$, and there are no single-qubit-gate and memory errors, which may be relevant for ion-trap [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67] and neutral-atom [68, 69, 70, 71, 72, 73, 74, 75, 76, 77] systems.
The logical-CNOT error probability is estimated as follows [104]. First, two error-free logical Bell pairs are prepared by performing error-free logical Hadamard and CNOT gates on error-free logical all-zero states. Next, we perform ten times transversal physical CNOT gates followed by error-correcting teleportation [16, 104, 116] on the first blocks of the two Bell pairs, where we use the encoders in Sec. IV and physical operations are faulty according to the above error model. Then, we disentangle the logical Bell pairs by performing error-free logical Hadamard and CNOT gates. Last, we measure all the physical qubits in the $Z$ basis and decode the results. We define the block error probability as $p_{\mathrm{block}}=1-(1-p_{10})^{1/10}$}, where $p_{10}$ is the probability that at least one of the logical-qubit values obtained by the decoding is 1. The error probability per logical CNOT, $p_{\mathrm{CNOT}}$, is also defined as $p_{\mathrm{CNOT}}=1-(1-p_{\mathrm{block}})^{1/K^{(3)}}$, where $K^{(3)}$ is the number of physical qubits in the level-3 block.
Figure 9 shows the results for the four level-3 MHC codes with the original and proposed encoders. All the exponents obtained by the power-function fitting are close to d/2=4{d/2=4} (d=8{d=8} is the code distance). This suggests the fault tolerance of the original and proposed encoders. Also, all the solid lines are a little lower than the corresponding dashed lines. This means that the proposed encoders are a little better than the original ones. Most importantly, $D_{6,4,4}$ achieve the lowest values for both $p_{\mathrm{block}}$ and $p_{\mathrm{CNOT}}$. This concludes that $D_{6,4,4}$ is the best of the level-3 MHC codes.
Figure 9: Numerical results of logical-CNOT error probability for level-3 MHC codes. (a–d) Log-log plots of the block error probability $p_{\mathrm{block}}$ for $D_{4,4,4}$, $D_{6,4,4}$, $D_{6,6,4}$, and $D_{6,6,6}$, respectively. $p_{\mathrm{block}}$ is the probability that at least one logical CNOT fails per block (see the main text for its definition). The circles and triangles represent the results with the proposed and original encoders, respectively. The solid and dashed lines are the fitting results with the power function $\beta p_{\mathrm{circ}}^{\alpha}$ ($\alpha$ and $\beta$ are fitting parameters) to the circles and triangles, respectively. In each panel, $p_{\mathrm{circ}}^{\alpha}$ shows the fitting result of $\alpha$. (e) $p_{\mathrm{block}}$ at $p_{\mathrm{circ}}=2\times 10^{-4}$}. (f) Logical-CNOT error probability $p_{\mathrm{CNOT}}$ at $p_{\mathrm{circ}}=2\times 10^{-4}$}. $p_{\mathrm{CNOT}}$ is defined as $p_{\mathrm{CNOT}}=1-(1-p_{\mathrm{block}})^{1/K^{(3)}}$, where $K^{(3)}$ is the number of physical qubits in the level-3 block.
Figure 10 shows the total number of physical qubits necessary for preparing a logial all-zero state, the increase due to the repetition when errors are detected is taken into account. The results clearly show that the proposed encoders can reduces the overheads compared with the original encoders even when the error-detection effect is included.
Figure 10: Total number of physical qubits necessary for preparing a logial all-zero state for level-3 MHC codes. (a–d) Semilog plots of the total number of physical qubits for $D_{4,4,4}$, $D_{6,4,4}$, $D_{6,6,4}$, and $D_{6,6,6}$, respectively. The solid and dashed curves show the results for the proposed and original encoders, respectively.
VI Summary and outlook
Toward lower logical error rates and earlier experimental realization of MHC codes, we have investigated the MHC codes with $[[6,4,2]]$ and/or $[[4,2,2]]$ quantum error-detecting codes. Consequently, we have found that $D_{6,4,4}$ and $D_{6,6,4,4}$ achieve the lowest decoding error probability among the level-3 and level-4 MHC codes, respectively. Focusing on level 3, we have also developed efficient fault-tolerant encoders achieving about 60% overhead reduction compared with the original encoders. The proposed encoders also improve the logical-CNOT performance a little. $D_{6,4,4}$ has also showed the highest performance for logical CNOT gates. Thus, we conclude that $D_{6,4,4}$ is the best of the level-3 MHC codes. This code can achieve not only lower logical error rates but also the smaller number of necessary physical qubits compared to the original $D_{6,6,6}$. Therefore, $D_{6,4,4}$ will become a good target toward experimental realization of FTQC with a high-rate quantum code.
In this work, we have evaluated only logical CNOT gates and left the evaluation of other logical gates for future work. Also, we have assumed no memory errors in the present circuit-level noise model. Toward experimental realization, e.g., with trapped ions or neutral atoms, we should optimize the gate sequence to minimize the moving distance of qubits. Such optimization is also left for future work.
Acknowledgements.
This work was supported by JST Moonshot R&D Grant Number JPMJMS2061. Parts of numerical simulations were done on the HOKUSAI supercomputer at RIKEN (project ID RB230022).